# NeuralFuse: Learning to Recover the Accuracy of Access-Limited Neural Network Inference in Low-Voltage Regimes

### v2

@@ Abstract
Deep neural networks (DNNs) have become ubiquitous in machine learning, but their energy consumption remains a notable issue. Lowering the supply voltage is an effective strategy for reducing energy consumption. However, aggressively scaling down the supply voltage can lead to accuracy degradation due to random bit flips in static random access memory (SRAM) where model parameters are stored. To address this challenge, we introduce NeuralFuse, a novel add-on module that addresses the accuracy-energy tradeoff in low-voltage regimes by learning input transformations to generate error-resistant data representations. NeuralFuse protects DNN accuracy in both nominal and low-voltage scenarios. Moreover, NeuralFuse is easy to implement and can be readily applied to DNNs with limited access, such as non-configurable hardware or remote access to cloud-based APIs. Experimental results demonstrate that, at a 1% bit error rate, NeuralFuse can reduce SRAM memory access energy by up to 24% while recovering accuracy by up to 57%. To the best of our knowledge, this is the first model-agnostic approach (i.e., no model retraining) to address low-voltage-induced bit errors. The source code is available at https://github.com/IBM/NeuralFuse.
@@ 1 Introduction
Energy-efficient computing is a primary consideration for the deployment of Deep Neural Networks (DNNs), particularly on edge devices and on-chip AI systems. Increasing energy efficiency and lowering the carbon footprint of DNN computation involves iterative efforts from both chip designers and algorithm developers. Processors with specialized hardware accelerators for AI computing are now ubiquitous, capable of providing orders of magnitude more performance and energy efficiency for AI computation. In addition to reduced precision/quantization and architectural optimizations, low voltage operation is a powerful knob that impacts power consumption. There is ample evidence in computer engineering literature that studies the effects of undervolting and low-voltage operation on accelerator memories that store weights and activations during computation. Aggressively scaling down the SRAM (Static Random Access Memory) supply voltage below the rated value leads to an exponential increase in bit failures, but saves power on account of the quadratic dependence of dynamic power on voltage (Chandramoorthy et al., 2019; Ganapathy et al., 2017). Such memory bit flips in the stored weight and activation values can cause catastrophic accuracy loss.
@@
A recent spate of works advocates low voltage operation of DNN accelerators using numerous techniques to preserve accuracy ranging from hardware-based error mitigation techniques (Chandramoorthy et al., 2019; Reagen et al., 2016) to error-aware robust training of DNN models (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021). On-chip error mitigation methods have significant performance and power overheads. On the other hand, some have also proposed to generate models that are more robust to bit errors via a specific learning algorithm (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021), thereby eliminating the need for on-chip error mitigation. However, error-aware training to find the optimal set of robust parameters for each model is time and energy-intensive and may not be possible in all access-limited settings.
@@
In this paper, we propose a novel model-agnostic approach: NeuralFuse. NeuralFuse is a machine learning module that allows for mitigating bit errors caused by very low voltage operation, through a trainable input transformation parameterized by a relatively small DNN, to enhance the robustness of the original input and provide accurate inference. The pipeline of NeuralFuse is illustrated in Figure 1. NeuralFuse accepts the scenarios under access-limited neural networks (e.g., non-configurable hardware or remote access to cloud-based APIs) to protect the deployed models from making wrong predictions under low power. Specifically, we consider two practical access-limited scenarios: (a) Relaxed Access, where the model details are unknown but backpropagation through the black-box models is possible, and (b) Restricted Access, where models are unknown, and backpropagation is disallowed. For relaxed access, we train NeuralFuse via backpropagation, and for restricted access, we train NeuralFuse on a white-box surrogate model and transfer it to the restricted access models. To the best of our knowledge, this is the first study that leverages a learning-based method to address random bit errors for recovering accuracy in low-voltage and access-limited settings.
@@
We propose NeuralFuse, a novel learning-based input transformation framework to enhance the accuracy of DNNs subject to random bit errors caused by undervolting. NeuralFuse is model-agnostic because it operates in a plug-and-play manner at the data input and does not require re-training the deployed DNN model.
We consider two practical access-limited scenarios for neural network inference: Relaxed Access and Restricted Access. In the former setting, we use gradient-based methods to train the NeuralFuse module. In the latter setting, we use a white-box surrogate model to train NeuralFuse and show its high transferability to other types of DNN architectures.
We conduct extensive experiments on various combinations of DNN models (ResNet18, ResNet50, VGG11, VGG16, and VGG19), datasets (CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10), and NeuralFuse implementations with different architectures and sizes. The results show that NeuralFuse can consistently increase the perturbed accuracy (accuracy evaluated under random bit errors in weights) by up to , while simultaneously saving the overall SRAM memory access energy by up to , based on the realistic characterization of bit cell failures for a given memory array in a low-voltage regime inducing a of bit error rate.
We demonstrate NeuralFuse’s advantages in various scenarios. The experimental results show that NeuralFuse has high transferability (adaptability to unseen base models), versatility (capable of recovering low-precision quantization loss, as well as bit errors), and competitiveness (achieving state-of-the-art performance).
@@ 2 Related Work and Background
Software-based Energy Saving Strategies.
@@
Recent studies have proposed reducing energy consumption from a software perspective. For instance, quantization techniques have been proposed to reduce the precision of storing model weights and decrease total memory storage (Gong et al., 2014; Rastegari et al., 2016; Wu et al., 2016). On the other hand, Yang et al. (2017) proposed energy-aware pruning on each layer and finetunes the weights to maximize the final accuracy. Yang et al. have also proposed several ways to reduce the energy consumption of DNN. For example, they proposed ECC, a DNN compression framework, that compresses a DNN model to meet the given energy constraint (Yang et al., 2019a); they have also proposed to compress DNN models via joint pruning and quantization (Yang et al., 2020). Besides, it is also feasible to treat energy constraint as an optimization problem during DNN training to reduce energy consumption and maximize training accuracy (Yang et al., 2019b). However, these methods focus on changing either the model architectures or model weights to reduce energy consumption, which is orthogonal to our approach with NeuralFuse, which serves as an add-on module for any given model.
Hardware-based Energy Saving Strategies.
@@
Existing works have studied improving energy efficiency by designing specific hardware. Several works have studied the undervolting of DNN accelerators and proposed methods to maintain accuracy in the presence of bit errors. For instance, Reagen et al. (2016) proposed an SRAM fault mitigation technique that rounds the faulty weights into zeros to avoid the degradation of the prediction accuracy. Srinivasan et al. (2016) proposed to store the sensitive MSBs (most significant bits) in robust SRAM cells to preserve accuracy. Chandramoorthy et al. (2019) proposed dynamic supply voltage boosting to a higher voltage to improve the resilience of the memory access operation. On the other hand, Stutz et al. (2021) considered a learning-based approach that tries to find models that are robust to bit errors. The paper discusses several techniques to improve the robustness, such as quantization methods, weight clipping, random bit error training, and adversarial bit error training. The authors observe that the combination of quantization, weight clipping, and adversarial bit error training achieves excellent performance in their experiments. However, they also admitted that the training process for that is sensitive to hyperparameter settings, and hence, it might come with a challenging training procedure.
@@
We argue that the methods mentioned above are not easy to implement or not suitable for real-world scenarios in access-limited settings. For example, the weights of DNN models packed on embedded systems may not be configurable or updatable. Therefore, model retraining (e.g., Stutz et al. (2021)) is not a viable option in this scenario. Moreover, the training of DNNs is already a tedious and time-consuming task. Adding error-aware training during training may further increase the training complexity and introduce challenges in hyperparameter search, as identified in previous literature. Özdenizci & Legenstein (2022) also note that this error-aware training was also found ineffective for large DNNs with millions of bits. Our proposed NerualFuse spares the need for model retraining by attaching a trainable input transformation function parameterized by a relatively small DNN as an add-on module to any DNN model as is.
@@
SRAM Bit Errors in DNNs. Low-voltage-induced memory bit cell failures can cause bit flips from 0 to 1 and vice versa. In practice, SRAM memory bit errors increase exponentially when the supply voltage is scaled below , which is the minimum voltage required to avoid bit errors. This phenomenon has been well studied in previous literature, such as the work by Chandramoorthy et al. (2019) and Ganapathy et al. (2017). As shown in Figure 2, for an SRAM array of size 51264 bits and 14nm technology node, the number of bit errors increases as the voltage scales down. The corresponding dynamic energy per read access of SRAM is shown on the right, measured at each voltage at a constant frequency. In this example, accessing the SRAM at 0.83 leads to a bit error rate, and at the same time, the dynamic energy per access is reduced by approximately . This can lead to inaccurate inferences in DNNs, particularly when bit flips occur at the MSBs. However, improving the robustness to bit errors allows us to lower the voltage and exploit the resulting energy savings.
@@
It has been observed that bit cell failures for a given memory array are randomly distributed and independent of each other. That is, the spatial distribution of bit flips could be assumed random, as it is generally different between arrays, even within a chip and between different chips. In this paper, we follow the methodology in Chandramoorthy et al. (2019) and model bit errors in a memory array of a given size by generating a random distribution of bit errors with equal likelihood of 0-to-1 and 1-to-0 bit flip. Specifically, we assume that the model weights are quantized to 8-bit precision (i.e., from 32-bit floats to 8-bit integers), and randomly distributed bit errors are injected into the quantized 2’s complement representation of weights to generate perturbed models. Please refer to Section 4.1 for more implementation details.
@@ 3.1 Error-Resistant Input Transformation
As illustrated in Figure 1, to overcome the drawback of performance degradation in low-voltage regimes, we propose a novel trainable input transformation module parametrized by a relatively small DNN, NeuralFuse, to mitigate the accuracy-energy tradeoff for model inference. The rationale is to use a specially designed loss function and training scheme to derive the NeuralFuse and apply it to the input data such that the transformed inputs will become robust to low-voltage-induced bit errors.
Consider the input sampled from the data distribution and a model with random bit errors on its weights (which is called a perturbed model). When there are no bit errors (i.e., the normal-voltage settings), the perturbed model reduces to a nominal deterministic model denoted by . NeuralFuse aims to ensure the perturbed model can make correct inferences on the transformed inputs as well as retain consistent results of in regular (normal-voltage) settings. In order to adapt to different data characteristics, NeuralFuse () is designed to be input-aware, which can be formally defined as:
where is a “generator” (i.e., an input transformation function) that can generate a perturbation based on the input . The NeuralFuse transformed data will be passed to the deployed model ( or ) for final inference. Without loss of generality, we assume the transformed input lies within a scaled input range , where is the (flattened) dimension of .
@@ 3.2 Training Objective and Optimizer
To train the generator , which should ensure the correctness of both the perturbed model and the clean model , we parameterize by a neural network and design the following training objective function:
where is the set of trainable parameters for , is the ground-truth label of , denotes the likelihood of computed by a model given a transformed input , is the distribution of the perturbed models inherited from the clean model under a random bit error rate, and is a hyperparameter that balances the importance between the nominal and perturbed models.
The training objective function can be readily converted to a loss function (loss) that evaluates the cross-entropy between the ground-truth label and the prediction . That is, the total loss function becomes
To optimize the loss function entailing the evaluation of the loss term on randomly perturbed models, our training process is inspired by the EOT (Expectation Over Transformation) attacks (Athalye et al., 2018). EOT attacks aim to find a robust adversarial example against a variety of image transformations. Based on the idea, we propose a new optimizer for solving (3), which we call Expectation Over Perturbed Models (EOPM). EOPM-trained generators can generate error-resistant input transformations and mitigate the inherent bit errors. However, it is computationally impossible to enumerate all possible perturbed models with random bit errors, and the number of realizations for perturbed models is limited by the memory constraint of GPUs used for training. In practice, we only take perturbed models for each iteration to calculate the empirical average loss, i.e.,
where is the number of simulated perturbed models under random bit errors to calculate the loss. Therefore, the gradient used to update the generator can be calculated as follows:
In our implementation, we find that can already deliver stable performance, and there is little gain in using a larger value. The ablation study for different can be found in Appendix E.
@@ 3.3 Training Algorithm
Algorithm 1 in Appendix A summarizes the training steps of NeuralFuse. We split the training data into mini-batches for training the generator in each epoch. For each mini-batch, we first feed these data into to get the transformed inputs. Also, we simulate perturbed models using a random bit error rate, denoted by , from . Then, the transformed inputs are fed into these perturbed models and the clean model to calculate their losses and gradients. Finally, NeuralFuse parameters are updated based on the gradient obtained by EOPM.

### v3

@@ Abstract
Deep neural networks (DNNs) have become ubiquitous in machine learning, but their energy consumption remains problematically high. An effective strategy for reducing such consumption is supply-voltage reduction, but if done too aggressively, it can lead to accuracy degradation. This is due to random bit-flips in static random access memory (SRAM), where model parameters are stored. To address this challenge, we have developed NeuralFuse, a novel add-on module that handles the energy-accuracy tradeoff in low-voltage regimes by learning input transformations and using them to generate error-resistant data representations, thereby protecting DNN accuracy in both nominal and low-voltage scenarios. As well as being easy to implement, NeuralFuse can be readily applied to DNNs with limited access, such cloud-based APIs that are accessed remotely or non-configurable hardware. Our experimental results demonstrate that, at a 1% bit-error rate, NeuralFuse can reduce SRAM access energy by up to 24% while recovering accuracy by up to 57%. To the best of our knowledge, this is the first approach to addressing low-voltage-induced bit errors that requires no model retraining.††footnotetext: Project Page: https://trustsafeai-neuralfuse.static.hf.space††footnotetext: Code: https://github.com/IBM/NeuralFuse
@@ 1 Introduction
Energy-efficient computing is of primary importance to the effective deployment of deep neural networks (DNNs), particularly in edge devices and in on-chip AI systems. Increasing DNN computation’s energy efficiency and lowering its carbon footprint require iterative efforts from both chip designers and algorithm developers. Processors with specialized hardware accelerators for AI computing, capable of providing orders of magnitude better performance and energy efficiency for AI computation, are now ubiquitous. However, alongside reduced precision/quantization and architectural optimizations, endowing such systems with the capacity for low-voltage operation is a powerful lever for reducing their power consumption.
The computer engineering literature contains ample evidence of the effects of undervolting and low-voltage operation on accelerator memories that store weights and activations during computation. Aggressive scaling-down of static random access memory’s (SRAM’s) supply voltage to below the rated value saves power, thanks to the quadratic dependence of dynamic power on voltage. Crucially, however, it also leads to an exponential increase in bit failures. Memory bit flips cause errors in the stored weight and activation values (Chandramoorthy et al., 2019; Ganapathy et al., 2017), leading to catastrophic accuracy loss.
@@
A recent wave of research has proposed numerous techniques for allowing low-voltage operation of DNN accelerators while preserving their accuracy. Most of these have been either hardware-based error-mitigation techniques or error-aware robust training of DNN models. On-chip error mitigation methods have significant performance and power overheads (Chandramoorthy et al., 2019; Reagen et al., 2016). On the other hand, some have proposed to generate models that are robust to bit errors via a specific learning algorithm (Kim et al., 2018; Koppula et al., 2019; Stutz et al., 2021), thereby eliminating the need for on-chip error mitigation. However, error-aware robust training to find the optimal set of robust parameters for each model is time- and energy-intensive and may not be possible in all access-limited settings.
@@
In this paper, therefore, we propose a novel model-agnostic approach: NeuralFuse. This proof-of-concept machine-learning module offers trainable input transformation parameterized by a relatively small DNN; and, by enhancing input’s robustness, it mitigates bit errors caused by very low-voltage operation, thus serving the wider goal of more accurate inferencing. The pipeline of NeuralFuse is illustrated in Figure 1. To protect the deployed models from making wrong predictions under low-power conditions, NeuralFuse accepts scenarios under access-limited neural networks (e.g., non-configurable hardware or cloud-based APIs). Specifically, we consider two access-limited scenarios that are common in the real world: 1) relaxed access, in which ‘black box’ model details are unknown, but backpropagation through those models is possible; and 2) restricted access, in which the model details are unknown and backpropagation is disallowed. To enable it to deal with relaxed access, we trained NeuralFuse via backpropagation, and for restricted-access cases, we trained it on a white-box surrogate model. To the best of our knowledge, this is the first study that leverages a learning-based method to address random bit errors as a means of recovering accuracy in low-voltage and access-limited settings.
@@
We propose NeuralFuse, a novel learning-based input-transformation module aimed at enhancing the accuracy of DNNs that are subject to random bit errors caused by very low voltage operation. NeuralFuse is model-agnostic, i.e., operates on a plug-and-play basis at the data-input stage and does not require any re-training of deployed DNN models.
We explore two practical limited-access scenarios for neural-network inference: relaxed access and restricted access. In the former setting, we use gradient-based methods for module training. In the latter one, we use a white-box surrogate model for training, which is highly transferable to other types of DNN architecture.
We report the results of an extensive program of experiments with various combinations of DNN models (ResNet18, ResNet50, VGG11, VGG16, and VGG19), datasets (CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10), and NeuralFuse implementations of different architectures and sizes. These show that NeuralFuse can consistently increase the perturbed accuracy (accuracy evaluated under random bit errors in weights) by up to , while simultaneously saving up to of the energy normally required for SRAM access, based on our realistic characterization of bit-cell failures for a given memory array in a low-voltage regime inducing a bit-error rate.
We demonstrate NeuralFuse’s transferability (i.e., adaptability to unseen base models), versatility (i.e., ability to recover low-precision quantization loss), and competitiveness (i.e., state-of-the-art performance) in various scenarios, establishing it as a promising proof-of-concept for energy-efficient, resilient DNN inference.
@@ 2 Related Work and Background
Software-based Energy-saving Strategies.
@@
Various recent studies have proposed software-based methods of reducing computing’s energy consumption. For instance, quantization techniques have been reported to reduce the precision required for storing model weights, and thus to decrease total memory storage (Gong et al., 2014; Rastegari et al., 2016; Wu et al., 2016). On the other hand, Yang et al. (2017) - who proposed energy-aware pruning on each layer and fine-tuning of weights to maximize final accuracy - suggested several ways to reduce DNNs’ energy consumption. For example, they devised the ECC framework, which compresses DNN models to meet a given energy constraint (Yang et al., 2019a), and a method of compressing such models via joint pruning and quantization (Yang et al., 2020). It is also feasible, during DNN training, to treat energy constraints as an optimization problem and thereby reduce energy consumption while maximizing training accuracy (Yang et al., 2019b). However, unlike ours, all these methods imply changing either model architectures or model weights.
Hardware-based Energy-saving Strategies.
@@
Prior studies have also explored ways of improving energy efficiency via specially designed hardware. Several of them have focused on the undervolting of DNN accelerators and proposed methods to maintain accuracy in the presence of bit errors. For instance, Reagen et al. (2016) proposed an SRAM fault-mitigation technique that rounds faulty weights to zero to avoid degradation of prediction accuracy. Srinivasan et al. (2016) recommended storing sensitive MSBs (most significant bits) in robust SRAM cells to preserve accuracy. Chandramoorthy et al. (2019) proposed dynamic supply-voltage boosting to improve the resilience of memory-access operations; and the learning-based approach proposed by Stutz et al. (2021) aims to find models that are robust to bit errors. The latter paper discusses several techniques for improving such robustness, notably quantization, weight-clipping, random bit-error training, and adversarial bit-error training. Its authors concluded from their experiments that a combination of quantization, weight-clipping, and adversarial bit-error training will yield excellent performance. However, they also admitted that the relevant training process was sensitive to hyperparameter settings, and hence, it might come with a challenging training procedure.
@@
However, we suggest that all the methods mentioned above are difficult to implement and/or unsuitable for use in real-world access-limited settings. For example, the weights of DNN models packed on embedded systems may not be configurable or updatable, making model retraining (e.g., Stutz et al. (2021)) non-viable in that scenario. Moreover, DNN training is already a tedious and time-consuming task, so adding error-aware training to it may further increase its complexity and, in particular, make hyperparameter searches more challenging. Özdenizci and Legenstein (2022) also reported that error-aware training was ineffective for large DNNs with millions of bits. NeuralFuse obviates the need for model retraining via an add-on trainable input-transformation function parameterized by a relatively small secondary DNN.
SRAM Bit Errors in DNNs.
@@
Low voltage-induced memory bit-cell failures can cause bit-flips from 0 to 1 and vice versa. In practice, SRAM bit errors increase exponentially when the supply voltage is scaled below , i.e., the minimum voltage required to avoid them. This phenomenon has been studied extensively in the prior literature, including work by Chandramoorthy et al. (2019) and Ganapathy et al. (2017). The specific increases in bit errors as voltage scales down, in the case of an SRAM array of bits with a 14nm technology node, is illustrated in Figure 2. The corresponding dynamic energy per SRAM read access, measured at each voltage at a constant frequency, is shown on the right-hand side of the figure. In this example, accessing the SRAM at 0.83 leads to a bit-error rate, and at the same time, dynamic energy per access is reduced by approximately . This can lead to DNNs making inaccurate inferences, particularly when bit-flips occur at the MSBs. However, improving robustness to bit errors can allow us to lower and exploit the resulting energy savings.
@@
It has been observed that bit-cell failures for a given memory array are randomly distributed and independent of each other. That is, the spatial distribution of bit-flips can be assumed to be random, as it generally differs from one array to another, within as well as between chips. Below, following Chandramoorthy et al. (2019), we model bit errors in a memory array of a given size by generating a random distribution of such errors with equal likelihood of 0-to-1 and 1-to-0 bit-flipping. More specifically, we assume that the model weights are quantized to 8-bit precision (i.e., from 32-bit floats to 8-bit integers), and generate perturbed models by injecting our randomly distributed bit errors into the two’s complement representation of weights. For more implementation details, please refer to Section 4.1.
@@ 3.1 Error-Resistant Input Transformation
As illustrated in Figure 1, we propose a novel trainable input-transformation module, NeuralFuse, parametrized by a relatively small DNN, to mitigate the accuracy-energy tradeoff for model inference and thus overcome the drawback of performance degradation in low-voltage regimes. A specially designed loss function and training scheme are used to derive NeuralFuse and apply it to the input data such that the transformed inputs will become robust to low voltage-induced bit errors.
Consider the input sampled from the data distribution and a set of models with random bit errors on weights (i.e., perturbed models). When it is not manifesting any bit errors (i.e., at normal-voltage settings), the perturbed model operates as a nominal deterministic one, denoted by . NeuralFuse aims to ensure that a model can make correct inferences on the transformed inputs while also delivering consistent results in its state. To adapt to various data characteristics, NeuralFuse – designated as in Eq. (1), below – is designed to be input-aware. This characteristic can be formally defined as
where is a “generator” (i.e., an input-transformation function) that can generate a perturbation based on input . As transformed by NeuralFuse, i.e., as , that input is passed to the deployed model ( or ) for final inference. Without loss of generality, we assume the transformed input lies within a scaled input range , where is the (flattened) dimension of .
@@ 3.2 Training Objective and Optimizer
To train our generator , which ought to be able to ensure the correctness of both the perturbed model and the clean model , we parameterized it with a neural network and apply our training objective function
where is the set of trainable parameters for ; is the ground-truth label of ; denotes the likelihood of as computed by a model being given a transformed input ; is the distribution of the perturbed models inherited from the clean model , under a random bit-error rate; and is a hyperparameter that balances the importance of the nominal and perturbed models.
The training objective function can be readily converted to a loss function () that evaluates cross-entropy between the ground-truth label and the prediction . That is, the total loss function can be calculated as
In particular, optimizing the loss function requires evaluation of the impact of the loss term on randomly perturbed models. Our training process is inspired by expectation over transformation (EOT) attacks (Athalye et al., 2018), which aim to produce robust adversarial examples that are simultaneously adversarial over the entire transformation distribution. Based on that idea, we propose a new optimizer for solving Eq. (3), which we call expectation over perturbed models (EOPM). EOPM-trained generators can generate error-resistant input transformations and mitigate inherent bit errors. However, it would be computationally impossible to enumerate all possible perturbed models with random bit errors, and the number of realizations for perturbed models is constrained by the memory size of the GPUs used for training. In practice, therefore, we only use perturbed models per iteration to calculate empirical average loss, i.e.,
where is the number of perturbed models that are simulated to calculate the loss caused by random bit errors. Therefore, the gradient used to update the generator can be calculated as follows:
Through our implementation, we found that stable performance could be delivered when , and that there was little to be gained by using a larger value. The results of our ablation study for different values of can be found in Appendix E.
@@ 3.3 Training Algorithm
Algorithm 1 in Appendix A summarizes NeuralFuse’s training steps. Briefly, this involves splitting the training data into mini-batches for training the generator in each epoch. For each mini-batch, we first feed these data into to obtain the transformed inputs. Also, we simulate perturbed models using a random bit-error rate, denoted by , from . Then, the transformed inputs are fed into those perturbed models as well as into the clean model , and their respective losses and gradients are calculated. Finally, NeuralFuse parameters are updated based on the gradient obtained by EOPM.